OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation x + 7y = -21. x intercept = (3, 0); y intercept = (0, 21) x intercept = (-3, 0); y intercept = (0, -21) x intercept = (21, 0); y intercept = (0, 3) x intercept = (-21, 0); y intercept = (0, -3)

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (midhun.madhu1987):

Write the above equation in the form y = mx + c, then c will be the y-intercept. Write the above equation in the form x = my + b, then b will be the x-intercept.

OpenStudy (anonymous):

write the equation in the form x/a+y/b=1 by dividing both sides by -21 from there a is the x-intercept and b is the y-intercept @kaylamarie081

OpenStudy (anonymous):

so the answer is option d